How to find flux through a segment which makes an angle theta at the center.

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Flux is the product if the electruc field and the surface area the charge covered.
Therfore, flux is given as:
Φ = ∫E⋅dA E = (kq)/r2 A = πa2 where a = r*sin(θ) 

E = Electric Field
R is the resultant vector 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)